OpenStudy (anonymous):

write the equation in slope-intercept form. what are the slope and y-intercept -3x - 10y = 7

OpenStudy (liamschumm):

Slope intercept form is y=mx+b.

OpenStudy (liamschumm):

So you need to put that in that form.

OpenStudy (liamschumm):

(I suggest first moving y to one side, and making the coefficient 1.)

OpenStudy (anonymous):

I took algebra 2 years ago, so could you walk me through this? I don't remember how to change equations like that :L

OpenStudy (liamschumm):

Ok, no problem.

OpenStudy (liamschumm):

So -3x-10y=7.

OpenStudy (liamschumm):

Then we can add 3x to both sides.

OpenStudy (liamschumm):

3x-3x-10y=7+3x.

OpenStudy (liamschumm):

10y=7+3x.

OpenStudy (liamschumm):

Ok?

OpenStudy (anonymous):

I get that you add them to both sides, yeah.

OpenStudy (liamschumm):

Now we divide both sides by 10.

OpenStudy (liamschumm):

\[y=\frac{7+3x}{10}.\]

OpenStudy (liamschumm):

Also, we know that \[\frac{a+b}{n}=\frac{a}{n}+\frac{b}{n}.\]

OpenStudy (liamschumm):

Do you see where to go next?

OpenStudy (anonymous):

Not at all... What are a and b?

OpenStudy (liamschumm):

That is an example. a and b can be anything.

OpenStudy (liamschumm):

So can n. (Except for 0.)

OpenStudy (liamschumm):

So\[\frac{1+1}{2}=\frac{1}{2}+\frac{1}{2}.\]

OpenStudy (liamschumm):

I am just showing that that rule works.

OpenStudy (anonymous):

Just so you see what I see

OpenStudy (anonymous):

but what do you plug in for n, a and b out of those numbers..?

OpenStudy (liamschumm):

y=(7+3x)/10.

OpenStudy (liamschumm):

So then...

OpenStudy (liamschumm):

\[\frac{7+3x}{10}=\frac{7}{10}+\frac{3x}{10}.\]

OpenStudy (liamschumm):

Therefore, we can write that as\[y=\frac{3}{10}x+\frac{7}{10}.\]

OpenStudy (liamschumm):

And in y=mx+b, m is the slope, and b is the y-intercept.

OpenStudy (anonymous):

So it's D?

OpenStudy (liamschumm):

Yes!

OpenStudy (anonymous):

Thanks :D

OpenStudy (liamschumm):

You are very welcome!